You’ve recently learned that the company where you work is being sold for $250,000. The company’s income statement indicates current profits of $12,000, which have yet to be paid out as dividends. Assuming the company will remain a “going concern” indefinitely and that the interest rate will remain constant at 9 percent, at what constant rate does the owner believe that profits will grow?

Instruction: Enter your response rounded to two decimal places.
Answer #1

Let g be the constant growth rate in profits

Value of the firm = $ 250,000

250000 = 12000*(1+g)/(9% -g)

22500 - 250000g = 12000 + 12000g

10500 = 262000 *g

g = 4.007= 4% approx